A)A mixture of xenon and helium gases, at a total pressure of 851 mm Hg, contains 30.4 grams of xenon and 0.519 grams of helium. What is the partial pressure of each gas in the mixture?

PXe =  mm Hg
PHe =  mm Hg

B)A mixture of helium and argon gases contains helium at a partial pressure of 487 mm Hg and argon at a partial pressure of 471 mm Hg. What is the mole fraction of each gas in the mixture?

XHe =
XAr =
Answer #1

According to Dalton's Law of partial pressure , Partial pressure of gas in mixture = X gas ( P total )

Where X gas is mole fraction of gas.

First calculate no. of moles & mole fraction of each gas.

We know that , No. of moles = Mass / Molar mass

\therefore No. of moles of Xe ( n Xe ) = 30.4 g / ( 131.30 g /mol ) = 0.2315 mol

\therefore No. of moles of He ( n He ) = 0.519 g / ( 4.00 g / mol ) = 0.1298 mol

Mole fraction of Xe = Moles of Xe / Total moles of gas = n Xe / ( n Xe + n He )

\therefore Mole fraction of Xe = 0.2315 / ( 0.2315 + 0.1298 ) = 0.641

Mole Fraction of He = n He / ( n Xe + n He )

\therefore Mole fraction of He = 0.1298 / ( 0.2315 + 0.1298 ) = 0.359

Partial pressure of Xe = 0.641 ( 851 mm Hg ) = 545 mm Hg

Partial pressure of He = 0.359 ( 851 mm Hg ) =306 mm Hg

ANSWER : P Xe = 545 mm Hg & P He = 306 mm Hg

PART 2

First calculate total pressure of gas

P total = P He + P A r

Therefore, P total = 487 mm Hg + 471 mm Hg = 958 mm Hg

According to Dalton's Law of partial pressure , Partial pressure of gas in mixture = X gas ( P total )

Where X gas is mole fraction of gas.

Therefore, X gas = Partial pressure of gas / P total

\therefore X He = 487 mm Hg / 958 mm Hg = 0.508

X Ar = 471 mm Hg / 958 mm Hg = 0.492

ANSWER : X He = 0.508 & X Ar = 0.492